# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

### v2

@@ Abstract
Large Language Models (LLMs)-based hybrid retrieval uses LLMs to encode queries and documents into low-dimensional dense or high-dimensional sparse vectors. It retrieves documents relevant to search queries based on vector similarities. Documents are pre-encoded offline, while queries arrive in real-time, necessitating an efficient online query encoder. Although LLMs significantly enhance retrieval capabilities, serving deeply parameterized LLMs slows down query inference throughput and increases demands for online deployment resources. In this paper, we propose LightRetriever, a novel LLM-based hybrid retriever with extremely lightweight query encoders. Our method retains a full-sized LLM for document encoding, but reduces the workload of query encoding to no more than an embedding lookup. Compared to serving a full-sized LLM on an H800 GPU, our approach achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse retrieval tasks, retaining an average of 95% full-sized performance.
@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [62]. The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval [29] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval [60] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 20] and BM25 [60]. Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [72] or ranks [15]. Text retrieval is broadly applied in various scenarios, such as web search [43], semantic textual similarity [8], fact checking [66], and retrieval augmented generation [39].
Figure 1: Architecture comparison of LLM-based hybrid retrieval with 1) Full-sized Query Inference (Previous works); and 2) Extreme Lightweight Query Inference (LightRetriever).
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [27, 60]. Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [79]. On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval [29, 11, 71] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers [71, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [37]. Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
@@
Modern Transformers-based LMs handle text modeling with the attention mechanism [68], including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts. As is the case for existing LLM-based text retrievers [71, 5], queries and documents are modeled equally with the same LLM, enduring the same computational burdens. In contrast, traditional statistical retrieval algorithms, such as the popular TF-IDF [63] and BM25 [60], only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics. They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference. Experiments from [65] show that BM25 still performs well in many retrieval tasks, even beating some previous BERT-based retrievers trained on MSMARCO [52].
@@
Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities. As is shown in Figure 1, we retain a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities, but completely remove the deep modeling of queries. 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder. The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts. All query token vectors are pre-computed and cached into a simple Embedding layer. Query inference costs are drastically reduced to only one Embedding lookup. 2) For sparse retrieval, similar to statistical retrieval algorithms, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency. is directly used as the sparse query vector. Thus, no more encoding stage for sparse retrieval is needed. The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity. These scores are linearly interpolated [72] for final hybrid retrieval.
@@
LightRetriever is super fast and lightweight with respect to query inference. Compared to serving a full-sized LLM (e.g., Mistral-7B [26]) on an H800 GPU, our method achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Mistral-7B, Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [65] and 8 Chinese tasks from CMTEB Retrieval [76]), retaining an average of 95% full-sized performance.
@@ 2 Algorithm
Figure 2: Full-sized symmetric dense retrieval for baseline training.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12]. The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [61], etc. Following [29], listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) [55], modern LM-based retrievers [64, 76, 70], especially LLM retrievers [71, 5, 38], customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although [64, 76] use instructions on both query and passage sides, most LLM-based retrievers [71, 5, 38] do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 [71] and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.4.
@@ 2.2 Dense retrieval of LightRetriever
Figure 3: Dense retrieval of LightRetriever.
@@
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 [71]. with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.5.,
@@
Following [9, 32], an auxiliary KL loss is also used to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (5)
Full-sized symmetric baselines could be trained with asymmetric encoders together. Thus, the computation of auxiliary loss is economical. Experiments show that this helps further narrow the performance gaps, but it is not mandatory.
@@ 2.3 Sparse retrieval of LightRetriever
Figure 4: Sparse retrieval of LightRetriever.
@@
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. Following SPLADE [20], to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
@@
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Following SPLADE [20], given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [56] is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@ 3.1 Settings
Datasets
@@
A total of 23 training sets (including 20 English and 3 Chinese) are used in our experiments by following [47, 71]. Most of them are available at preprocessed data  [^4] https://huggingface.co/datasets/tdro-llm/finetune_data from tDRO [47] and Sentence Transformers Training Data  [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data [59]. Detailed training sets and sampling ratios are listed in the Appendix A.3. BeIR [65] and CMTEB Retrieval [76] benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following [65, 76], accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Models & hyper-parameters
@@
We modify the experiment code based on the implementation  [^6] https://github.com/ma787639046/tdro from tDRO [47]. All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE [20], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to maximum at first 4k steps, which helps reduce the side effects of the regulator at initial steps [56]. LoRA [24] is used to save GPU memory usage during the LLM training, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Mistral-v0.3-7B [26], Llama-3.2-1B, 3B, Llama-3.1-8B [67], and Qwen-2.5-1.5B, 3B, 7B [2]. Direct linear interpolation [72] is used to combine dense and sparse similarity scores as the hybrid retrieval scores. All results are conducted on a DGX node with 8 NVIDIA H800 GPUs, 128 CPU cores of 2 Intel Xeon 8462Y+, and 2 TB of memory.
Baselines
Full-sized symmetric dense retrievers are our main baselines under the same training conditions, as shown in Figure 2. The query inference speeds and retrieval performance gaps are compared as the main results. Additionally, we also include some recent retrievers in the results, including Sentence Transformers static embedding [1], BM25 [60], BGE-m3 [11], E5-Mistral [71], and LLM2Vec [5]. However, they are not our main comparison baselines, because [1] and [60] are parameter-efficient but less performant. LLM-based retrievers [71, 5] are performant but inefficient.
Table 1: Comparison of query encoding time and speedup for full-sized baseline and LightRetriever. Tokenization time is also presented for reference. 65536 queries from MSMARCO [52] are processed. Evaluations are conducted on an NVIDIA DGX node using 1 H800 GPU with Mistral-v0.3-7B.
Tokenzation Baseline LightRetriever
GPU Encode No-GPU Encode GPU Encode
Batch Size Time (s) Time (s) Time (s) speedup Time (s) speedup
16 5.1168 69.9276 3.7065 19 0.1509 463
64 3.8995 42.2991 3.1719 13 0.0510 829
256 2.0713 45.8708 2.9990 15 0.0185 2480
1024 1.4226 57.4545 2.7969 21 0.0107 5370
4096 1.6798 74.7221 2.8218 26 0.0093 8035
8192 OOM 2.8529 0.0090
3.2 Main results
Encoding speeds
The encoding speed comparisons of full-sized baseline and LightRetriever are presented in Table 1. No-GPU and GPU settings are measured for LightRetriever. Because of the efficiency issue, the baseline is only measured on the GPU. As is shown in Table 1, it takes a full-sized Mistral-7B LLM up to 74s for encoding 65536 queries on an H800 GPU across different batch sizes. In comparison, our LightRetriever only consumes less than 0.1509s on the GPU, which achieves over 400-8000x speedup. Even when encoding without GPU acceleration, LightRetriever also achieves over 10-20x speedup compared to the GPU baseline.
@@ 3.2 Main results
Table 2: Retrieval performances of LightRetriever on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks. The nDCG@10 gaps between LightRetriever hybrid and the corresponding baseline are presented. The best metrics of baselines and LightRetrievers are marked in bold.

### v3

@@ Abstract
Large Language Models (LLMs)-based text retrieval retrieves documents relevant to search queries based on vector similarities. Documents are pre-encoded offline, while queries arrive in real-time, necessitating an efficient online query encoder. Although LLMs significantly enhance retrieval capabilities, serving deeply parameterized LLMs slows down query inference throughput and increases demands for online deployment resources. In this paper, we propose LightRetriever, a novel LLM-based retriever with extremely lightweight query encoders. Our method retains a full-sized LLM for document encoding, but reduces the workload of query encoding to no more than an embedding lookup. Compared to serving a full LLM on an A800 GPU, our method achieves over 1000x speedup in query encoding and over 10× increase in end-to-end retrieval throughput. Extensive experiments on large-scale retrieval benchmarks show that LightRetriever generalizes well across diverse tasks, maintaining an average of 95% retrieval performance.
@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities Salton (1975). The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval Karpukhin et al. (2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval Robertson et al. (1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) Bai et al. (2020); Formal et al. (2021) and BM25 Robertson et al. (1994). Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores Wang et al. (2021) or ranks Cormack et al. (2009). Text retrieval is broadly applied in various scenarios, such as web search Lu et al. (2022), semantic textual similarity Bowman et al. (2015), fact checking Thorne et al. (2018), and retrieval augmented generation Lewis et al. (2020).
Figure 1: LightRetriever targets at extreme query inference speed for LLM-based text retrieval, reducing the workload of query encoding to no more than an embedding lookup.
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases Johnson et al. (2019); Robertson et al. (1994). Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere Yang et al. (2018a). On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval Karpukhin et al. (2020); Chen et al. (2024); Wang et al. (2024b) utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers Wang et al. (2024b); BehnamGhader et al. (2024) employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources Kwon et al. (2023). Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
@@
Modern Transformers-based LMs handle text modeling with the attention mechanism Vaswani et al. (2017), including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts. As is the case for existing LLM-based text retrievers Wang et al. (2024b); BehnamGhader et al. (2024), queries and documents are modeled equally with the same LLM, enduring the same computational burdens. In contrast, traditional statistical retrieval algorithms, such as the popular BM25 Robertson et al. (1994), only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics. They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference. Experiments from Thakur et al. (2021) show that BM25 still performs well in many retrieval tasks.
@@
Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities. Our work retains a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities. But as is shown in Figure 1, we completely remove the deep modeling of queries. 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder. The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts. All query token vectors are pre-computed and cached into a simple Embedding layer. Query inference costs are drastically reduced to only one Embedding lookup. 2) For sparse retrieval, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency. is directly used as the sparse query vector. Thus, no more encoding stage for sparse retrieval is needed. The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity. These scores are summed Wang et al. (2021) for the final hybrid retrieval.
@@
LightRetriever is super fast and lightweight for query inference. Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR Thakur et al. (2021) and 8 Chinese tasks from CMTEB Retrieval Xiao et al. (2023)), maintaining an average of 95% retrieval performance.
@@ 2 Algorithm
Figure 2: Contrastive training of full-sized symmetric dense retrieval. A full-sized dual-encoder is used to model both queries and documents. Task-specific instructions are added as common practices Su et al. (2023); Wang et al. (2024b) to promote better domain adaptation abilities.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ Chen et al. (2009). The above objective can be effectively optimized by the pairwise/listwise contrastive loss Cao et al. (2007), margin loss Rosset et al. (2003), etc. Following Karpukhin et al. (2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) Ouyang et al. (2022), modern LM-based retrievers Su et al. (2023); Xiao et al. (2023); Wang et al. (2024a), especially LLM retrievers Wang et al. (2024b); BehnamGhader et al. (2024); Lee et al. (2024), customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although Su et al. (2023); Xiao et al. (2023) use instructions on both query and passage sides, most LLM-based retrievers Wang et al. (2024b); BehnamGhader et al. (2024); Lee et al. (2024) do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 Wang et al. (2024b) and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.6.
@@ 2.2 Dense retrieval of LightRetriever
Figure 3: Dense retrieval of LightRetriever. Three stages are involved for efficient query modeling. 1) Training: A shared prompt concatenated with a single query token is passed through a full encoder independently to obtain token-level representations. The full query representation is then computed by averaging the token vectors corresponding to each query token. 2) Caching: Prior to serving, the trained encoder is used to precompute token embeddings for the entire LLM’s vocabulary, which are stored in a single embedding lookup matrix. 3) Online Serving: At inference time, query embeddings are efficiently generated by lookup and averaging the cached token embeddings, eliminating the need for deep model inference.
@@
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 Wang et al. (2024b). with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
@@ 2.3 Sparse retrieval of LightRetriever
Figure 4: Sparse retrieval of LightRetriever. The query representation is a term-based, unlearnable sparse vector, which is obtained from tokenization. The document representation is end-to-end trained by projecting last layer hidden states of LLM to vocabulary space, then optimized via contrastive learning.
@@
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. Following SPLADE Formal et al. (2021), to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
@@
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Following SPLADE Formal et al. (2021), given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator Paria et al. (2020) is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
Table 1: Comparison of retrieval speed and effectiveness. We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset Nguyen et al. (2016), evaluating both full symmetric retrievers and the query-lightweight LightRetriever. To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time. The overall throughput is measured in Queries Per Second (QPS).
Benchmark / nDCG@10 Time consumption / s & Throughput / QPS
Model BeIR CMTEB-R Tokenize Encode Search Total QPS
Full-Llama8b 56.8 67.6 1.3746 109.4853 8.5133 119.3730 549
Full-Llama1b 53.1 63.8 1.2564 19.9030 8.4897 29.6490 2210
LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412 8.5010 9.3630 6999
Full-Qwen7b 56.6 70.1 1.4144 100.6716 8.5279 110.6140 592
Full-Qwen1.5b 54.0 65.9 1.4210 53.0858 8.5299 63.0370 1040
LightRetriever-Qwen7b 53.8 66.5 0.8300 0.0420 8.5150 9.3870 6982
@@ 3.1 Settings
Training data
Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity. To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO Ma et al. (2025) and Sentence Transformers Training Data Reimers (2019). English datasets include Amazon Review Ni et al. (2019), Eli5 Fan et al. (2019), FEVER Thorne et al. (2018), FiQA Maia et al. (2018), GooQA Khashabi et al. (2021), HotpotQA Yang et al. (2018b), MSMARCO Nguyen et al. (2016), NFCorpus Boteva et al. (2016), NPR Lucas et al. (2023), NQ Kwiatkowski et al. (2019), PAQ Lewis et al. (2021), Quora Duplicates Iyer et al. (2012), S2ORC Lo et al. (2020), Scifact Wadden et al. (2020), Specter Cohan et al. (2020), StackExchange Duplicates da Silva et al. (2018), E5 synthetic BeastyZ (2024), Trivia Joshi et al. (2017), WikiHow Koupaee and Wang (2018), and Yahoo Answers Zhang et al. (2015). Chinese datasets include cMedQA2 Zhang et al. (2018), DuReader Qiu et al. (2022), and T2Ranking Xie et al. (2023).
Benchmarks
@@
BeIR Thakur et al. (2021) and CMTEB Retrieval Xiao et al. (2023) benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following Thakur et al. (2021); Xiao et al. (2023), accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Training hyper-parameters
@@
Our experiments are conducted based on the implementation of tDRO Ma et al. (2025). All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE Formal et al. (2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps Paria et al. (2020). LoRA Hu et al. (2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B Touvron et al. (2023), and Qwen-2.5-1.5B, 3B, 7B Bai et al. (2023). The hybrid similarity scores are linearly summed from normalized dense and sparse scores Wang et al. (2021). All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
Speed comparisons
To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCONguyen et al. (2016). All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search. Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
@@
Table 2: Performance comparisons on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks. The best metrics of full symmetric retrievers and LightRetrievers are marked in bold. Their nDCG@10 gaps are also presented.